# Filling in the Blanks:

### v1

@@ 5 Testing $v_{\mathrm{los}}$ Predictions
$H_{\mathrm{True}}=H\left(v_{\mathrm{True}}|\bm{u}\right)=\int_{-\infty}^{v_{\mathrm{True}}}p\left(v|\bm{u}\right)\mathrm{d}v$ (18)
@@
$\Delta_{\%}\approx\frac{1}{2}\sum_{i}^{\mathrm{bins}}\left|f^{\mathrm{Hist}}\left(H_{i}\right)-1\right|\Delta_{x}$ (19)
@@
$\Delta_{\mathrm{KS}}=100\times\max\left|H^{\mathrm{Hist}}\left(H_{\mathrm{True}}\right)-H^{\mathrm{Uniform}}\left(H_{\mathrm{True}}\right)\right|$ (20)
@@
5.1 Constraining $v_{\mathrm{los}}$
Figure 6: The distribution of $\sigma_{\mathrm{PDF}}$ (defined as half the $84\%-16\%$ $v_{\mathrm{los}}$ values) of the PDFs predicted by our method. The sample is split into Halo and Disc stars, whose normalised histograms of the true decomposition are plotted with solid-filled colours. The likely decomposition, defined by the probability of membership greater than 80%, is plotted with lines. The median values for the distributions for stars belonging to different groups (based on Dodd23) are shown in different colours.
@@ 5.1 Constraining $v_{\mathrm{los}}$
$\sigma_{\mathrm{PDF}}=\frac{1}{2}\left(v_{84\%}-v_{16\%}\right)\$ (21)
@@
5.2 Predicting $v_{\mathrm{los}}$
Figure 7: The distribution of the difference between $v_{\mathrm{los}}$ predictions and the true $v_{\mathrm{los}}$ across different subsamples of the Disc, the total Halo, and the stars belonging to substructure in the Halo. The top panel sample is from the Gaia RVS catalogue and includes $v_{\mathrm{los}}$ estimates our PDF’s the median and peak maximum values. The bottom panel shows a sample of stars from surveys crossmatched with Gaia and includes $v_{\mathrm{los}}$ estimates from Naik24. The tables contain the median velocity error for the different velocity predictions across the different subsamples.
Figure 8: The uncertainty in our predicted $v_{\mathrm{los}}$ PDFs against the difference between the true value $v_{\mathrm{True}}$ and our predicted line-of-sight velocity $v_{\mathrm{max}}$. True Disc stars are plotted in purple, downsampled by a factor of 1000, the true Halo stars that are determined to be likely Halo stars by our method (probability greater than $80\%$) in orange, and true Halo stars that are misdetermined to be likely disc stars in red, both downsampled by a factor of 2. The median line and shaded regions indicate the binned running median and $1\sigma$ and $2\sigma$ values of the total distribution, defined from the $16\%-84\%$ range. If the predictions are unbiased and the predicted PDF uncertainties accurate, the binned $1\sigma$ and $2\sigma$ values regions are expected to follow the value of their own corresponding bin, plotted in solid black and blue lines.
@@ 5.2 Predicting $v_{\mathrm{los}}$
5.3 Marginalising over $v_{\mathrm{los}}$
@@
Figure 10: The cylindrical velocity distributions of the different groups, as recovered by our methodology (lefthand side) against the true distributions (righthand side). Different stellar groups are represented by different colours, with the opacity of the colour representing the $\left(0.5\sigma,1\sigma,2\sigma\right)$ levels (equivalently the $38\%$, $64\%$, and $95\%$ regions). The groups are split into two separate subpanels, with the majority of phase-mixed groups in the top set and phase-clumped groups, with the addition of the Helmi Streams to allow the groups to be visually distinguished, in the bottom set.
@@ 6 Predicting Membership Probability
Figure 11: The distribution of membership probabilities of our individual groups (separate panels), calculated with our methodology. The distribution of the membership probabilities of every star is plotted in blue, and the true group members in orange. The first two panels depict the Disc and the total Halo. The total Halo is decomposed into individual structures and Ungrouped Halo. By construction, a star’s probabilities of belonging to the Disc or the total Halo sum to one, and the total Halo membership is then decomposed into belonging to structures or Ungrouped Halo.
@@ 6.1 Expected Populations
Figure 12: The expected populations of the stellar groups of our 6D RVS sample, calculated using membership probabilities assigned by our method, compared to the true population (based on Dodd23 groups). The error bars are estimated by redrawing the populations based on the membership probabilities of the stars.
@@
$N_{g}\right>=\sum_{i\in\mathrm{Stars}}p_{g}^{i}$ (22)
@@ 6.2 Purity, Completeness and Selecting a Sample
Figure 13: For each group (panel), we consider the effects of choosing a minimum membership probability to select a group of stars belonging to each group. Increasing the probability cut on membership probability (x-axis) increases the purity of the selected group, but decreases the completeness (y-axis). We plot the true purity and completeness, and the expected purity and completeness calculated from the probability distribution. For groups that show evidence of being phase-clumped, we also plot the purity and completeness calculated without the additional phase considerations (see Sec. 4.2).
@@
$C_{g}\left(p_{\min}\right)$ $=N_{g}^{\cap}\left(p_{\min}\right)/N_{\mathrm{g}}$ (24)
@@
$N_{g}\right>\left(p_{\min}\right)=\sum_{i\in\left[p_{g}^{i}>p_{\mathrm{cut}}\right]}p_{g}^{i}$ (25)
@@ 7.1 Comparison to Naik24
Figure 15: A comparison between the differences between the true $v_{\mathrm{los}}$ and the predictions of our methodology and Naik24, with samples decomposed into different halo substructures. We include all stars from the true groups, plotted with triangles, and true likely members with membership probability greater than $50\%$. Only groups with more than 4 likely members are plotted. The points and error bars represent the median and the $16\%$ and $84\%$ percentiles of the distribution, respectively, for the likely groups. These results use the crossmatched 6D sample described in the main text.
@@ 7.2 Comparison to Mikkola23
In the Halo, their predicted velocity distributions show the signature of several prominent accreted structures, which also appear in their recovered action space distributions. By finding the total density within these velocity features, they estimate the population of substructures of the 5D sample. In theory, the methodology could be extended to attribute individual stars to particular structures, but without this step, we cannot currently directly compare predictions for single stars.
7.3 Selecting a Halo Sample with $V_{\mathrm{Tan}}$
@@
$V_{\mathrm{Tan}}=4.74057\,\mathrm{km}\mathrm{s}^{-1}\left(\frac{\mu}{\mathrm{mas}}\mathrm{yr}^{-1}\right)\left(\frac{d}{\mathrm{kpc}}\right)$ (26)
@@
Any star with a large $V_{\mathrm{Tan}}$ is likely to have a large $V_{\mathrm{Toomre}}$ and can be assumed to be kinematically a Halo star. However, the selection on $V_{\mathrm{Tan}}$ neglects Halo stars that distinguish themselves from the Disc through high $v_{\mathrm{los}}$. Therefore, a conservative cut in $V_{\mathrm{Tan}}$ gives a relatively pure but incomplete selection in $V_{\mathrm{Toomre}}$. The halo selection cut typically used is $V_{\mathrm{Tan}}>200\,\text{km}\hskip 1.00006pt\text{s}^{-1}$ (Gaia Collaboration et al., 2018b; Mikkola et al., 2023), which selects a $V_{\mathrm{Toomre}}$ defined Halo with $83\%$ purity and $68\%$ completeness. A stricter cut improves purity but quickly reduces the completeness, as shown in the middle panel of Fig. 16. $V_{\mathrm{Tan}}>300\,\text{km}\hskip 1.00006pt\text{s}^{-1}$ gives a sample of greater than $99\%$ purity but only $21\%$ completeness.
@@ 8.2 Improving the Method
Our methodology could be developed further in several ways. First, we have restricted ourselves to a sample within $2.5\,\mathrm{kpc}$ to ensure a high-quality, well-sampled representation of the phase-space distribution. If we could extend our methodology to greater distances, we would have a greater number of stars and additional structures not seen within the local solar neighbourhood, for example, the Sagittarius stream. This extension would require a sufficiently large 6D sample of stars at larger distances to model the phase space effectively. Alternatively, we could model the action distributions of the Disc and Halo using analytic distributions (li21ModellingStellarHalo; Binney & Vasiliev, 2024). These distributions would allow us to find the likely Disc and Halo membership probability for 5D stars at any distance and not suffer from poor sampling in high-energy regions. However, the Halo is arguably composed entirely of substructure (Naidu et al., 2020), the majority material of GES. Any functional distribution would be imperfect and could easily dominate our smaller substructures or cause biases in poorly fit regions of the phase space. Instead, we have chosen to remain as close to the data as possible.
@@ 1. Observational Probability of a Single Star at a Single Point
First, we must be able to estimate the probability that a given star, with known stellar properties and position is in our data sample. When a star of intrinsic absolute magnitude and colour is observed, its apparent magnitude and colour are affected by extinction, depending on the local dust distribution and its physical position relative to the Sun. We use the dust map of Lallement et al. (2022) and the coefficients of Gaia Collaboration et al. (2018b) to apply reddening and de-reddening corrections.
The Gaia selection function is complex and the subject of various studies (Boubert & Everall, 2022; Castro-Ginard et al., 2023). In Castro-Ginard et al. (2023), the RVS selection function has been estimated to depend on the sky position and apparent magnitude and colour of the star $\left(l,b,G,G-G_{RP}\right)$. Given these observables, these authors provide a method to estimate the probability that the star would be included in the Gaia RVS catalogue. Numerically, the RVS function information is defined with a physical grid of healpix bins and distances, followed by a grid of colours and magnitudes.
@@ 2. Expected Observed Fraction of a Halo Population at a Single Point
Instead of considering the probability of seeing a single star at a point in space, we must consider the expected fraction observed of a typical Halo stellar population. To do this, we draw a stellar population from an isochrone, using a Chabrier (Chabrier, 2003) Initial Mass Function (IMF). We use an alpha-enhanced BaSTI isochrone (Pietrinferni et al., 2021) with a metallicity of $\left[\mathrm{Fe}/\mathrm{H}\right]=-1.1$ and an age of $11$ Gyrs.
@@ Appendix B Density Estimation
$\rho\left(\sqrt{J_{R}+\delta_{J}},\sqrt{J_{z}+\delta_{J}},L_{z}\right)$
@@
$\rho\left(\tilde{\bm{J}}|\left[\tilde{\bm{J}}_{i},w_{i}\right]\right)={\left(\sigma_{R}\sigma_{Z}\sigma_{\phi}\right)}^{-1}\sum_{i=1}^{N}w_{i}{\lambda_{i}}^{-3}K_{e}\left(\frac{1}{\lambda_{i}}\left|\frac{\tilde{\bm{J}}-\tilde{\bm{J}}_{i}}{\bm{\sigma}}\right|\right)$ (27)
@@
$K_{e}\left(r\right)=\begin{cases}\frac{d+2}{2V_{d}}\left(1-r^{2}\right)&r^{2}<1\\ 0&\text{otherwise}\end{cases}$ (28)
@@
$\sigma_{j}=\frac{P_{100-\omega\%}\left(\tilde{J}_{j}\right)-P_{\omega\%}\left(\tilde{J}_{j}\right)}{\log\left(N\right)}$ (29)
@@
$\lambda_{i}={\left(\frac{\hat{\rho}\left(\tilde{\bm{J}}_{i}\right)}{g}\right)}^{-\frac{1}{3}}$ (30)
@@ Appendix C Disc Disequillibirum
Instead, we look at the dynamics within the Disc using the cylindrical components of velocity $\left(v_{R},v_{\phi},v_{Z}\right)$. We consider the deviation from expected equilibrium values to quantify disequilibrium in velocity. If the Disc is in phase-mixed equilibrium in every bin, we expect even numbers of stars moving up and down through the disc and in and out towards the centre of the Galaxy. Defining $N_{+}$ and $N_{-}$ as the number of stars with a positive and negative velocity sign, respectively, we can estimate the statistical significance of any imbalance by modelling the distribution as a binomial, which can be approximated by a normal distribution. The statistical likelihood of getting $\left(N_{+},N_{-}\right)$ in units of sigma is then $\chi$:
$\chi=\frac{1}{\sqrt{N}}\left(N_{+}-N_{-}\right)$ (31)
@@
We plot these $\chi$ values for each of the velocity components in the bottom panels. They show strong, clear evidence for dynamical disequilibrium in Disc stars. The left panel showing the tangential value has very good agreement with the local spiral arm structure, as can be seen in Poggio et al. (2021) Fig. 5. The middle radial panel is consistent with Vislosky et al. (2024), and at the solar position, we are at the edge of a structure that is moving radially outward. The final panel of vertical motion shows a slight structure at a larger radius, away from the centre. There is more structure when comparing above and below the disc plane, which is likely linked to the Gaia phase space spiral (Antoja et al., 2018; Hunt et al., 2022).

### v2

@@ 5 Testing vlos predictions
$H_{\mathrm{True}}=H\left(v_{\mathrm{True}}|\bm{u}\right)=\int_{-\infty}^{v_{\mathrm{True}}}p\left(v|\bm{u}\right)\mathrm{d}v.$ (18)
@@
$\Delta_{\%}\approx\frac{1}{2}\sum_{i}^{\mathrm{bins}}\left|f^{\mathrm{Hist}}\left(H_{i}\right)-1\right|\Delta_{x},$ (19)
@@
$\Delta_{\mathrm{KS}}=100\times\max\left|H^{\mathrm{Hist}}\left(H_{\mathrm{True}}\right)-H^{\mathrm{Uniform}}\left(H_{\mathrm{True}}\right)\right|.$ (20)
@@
5.1 Constraining vlos
Figure 6: Distributions of $\sigma_{\mathrm{PDF}}$ (defined as half the $84\%-16\%$ $v_{\mathrm{los}}$ values) of the PDFs predicted by our method. The sample is split into Halo and Disc stars, whose normalised histograms of the true decomposition are plotted with solid-filled colours. The likely decomposition, defined by the probability of membership greater than 80%, is plotted with lines. The median values for the distributions for stars belonging to different groups (based on Dodd23) are shown in different colours.
@@ 5.1 Constraining vlos
$\sigma_{\mathrm{PDF}}=\frac{1}{2}\left(v_{84\%}-v_{16\%}\right),$ (21)
@@
5.2 Predicting vlos
Figure 7: Distributions of the difference between $v_{\mathrm{los}}$ predictions and the true $v_{\mathrm{los}}$ across different subsamples of the Disc, the total Halo, and the stars belonging to substructure in the Halo. The top panel sample is from the Gaia RVS catalogue and includes $v_{\mathrm{los}}$ estimates our PDF’s the median and peak maximum values. The bottom panel shows a sample of stars from surveys crossmatched with Gaia and includes $v_{\mathrm{los}}$ estimates from Naik24. The tables contain the median velocity error for the different velocity predictions across the different subsamples.
Figure 8: Uncertainty in our predicted $v_{\mathrm{los}}$ PDFs against the difference between the true value $v_{\mathrm{True}}$ and our predicted line-of-sight velocity $v_{\mathrm{max}}$. True Disc stars are plotted in purple, downsampled by a factor of 1000, the true Halo stars that are determined to be likely Halo stars by our method (probability greater than $80\%$) in orange, and true Halo stars that are misdetermined to be likely disc stars in red, both downsampled by a factor of 2. The median line and shaded regions indicate the binned running median and $1\sigma$ and $2\sigma$ values of the total distribution, defined from the $16\%-84\%$ range. If the predictions are unbiased and the predicted PDF uncertainties accurate, the binned $1\sigma$ and $2\sigma$ values regions are expected to follow the value of their own corresponding bin, plotted in solid black and blue lines.
@@ 5.2 Predicting vlos
5.3 Marginalising over vlos
@@
Figure 10: Cylindrical velocity distributions of the different groups, as recovered by our methodology (lefthand side) against the true distributions (righthand side). Different stellar groups are represented by different colours, with the opacity of the colour representing the $\left(0.5\sigma,1\sigma,2\sigma\right)$ levels (equivalently the $38\%$, $64\%$, and $95\%$ regions). The groups are split into two separate subpanels, with the majority of phase-mixed groups in the top set and phase-clumped groups, with the addition of the Helmi Streams to allow the groups to be visually distinguished, in the bottom set.
@@ 6 Predicting membership probability
Figure 11: Distributions of membership probabilities of our individual groups (separate panels), calculated with our methodology. The distribution of the membership probabilities of every star is plotted in blue, and the true group members in orange. The first two panels depict the Disc and the total Halo. The total Halo is decomposed into individual structures and Ungrouped Halo. By construction, a star’s probabilities of belonging to the Disc or the total Halo sum to one, and the total Halo membership is then decomposed into belonging to structures or Ungrouped Halo.
@@ 6.1 Expected populations
Figure 12: Expected populations of the stellar groups of our 6D RVS sample, calculated using membership probabilities assigned by our method, compared to the true population (based on Dodd23 groups). The error bars are estimated by redrawing the populations based on the membership probabilities of the stars.
@@
$N_{g}\right>=\sum_{i\in\mathrm{Stars}}p_{g}^{i},$ (22)
@@ 6.2 Purity, completeness and selecting a sample
Figure 13: Plots depicting effects of increasing the minimum membership probability upon the purity and completeness of the selected sample of a group. For each group (panel), increasing the probability cut on membership probability (x-axis) increases the purity of the selected group, but decreases the completeness (y-axis). We plot the true purity and completeness, and the expected purity and completeness calculated from the probability distribution. For groups that show evidence of being phase-clumped, we also plot the purity and completeness calculated without the additional phase considerations (see Sec. 4.2).
@@
$C_{g}\left(p_{\min}\right)$ $=N_{g}^{\cap}\left(p_{\min}\right)/N_{\mathrm{g}},$ (24)
@@
$N_{g}\right>\left(p_{\min}\right)=\sum_{i\in\left[p_{g}^{i}>p_{\mathrm{cut}}\right]}p_{g}^{i}.$ (25)
@@ 7.1 Comparison to Naik24
Figure 15: Comparisons between the differences between the true $v_{\mathrm{los}}$ and the predictions of our methodology and Naik24, with samples decomposed into different halo substructures. We include all stars from the true groups, plotted with triangles, and true likely members with membership probability greater than $50\%$. Only groups with more than 4 likely members are plotted. The points and error bars represent the median and the $16\%$ and $84\%$ percentiles of the distribution, respectively, for the likely groups. These results use the crossmatched 6D sample described in the main text.
@@ 7.2 Comparison to Mikkola23
In the Halo, their predicted velocity distributions show the signature of several prominent accreted structures, which also appear in their recovered action space distributions. By finding the total density within these velocity features, they estimate the population of substructures of the 5D sample. In theory, the methodology can also be extended to attribute individual stars to particular structures (e.g., Cronstedt, 2023) but we cannot currently directly compare predictions for single stars.
7.3 Selecting a halo sample with vtan
@@
$V_{\mathrm{Tan}}=4.74057\,\mathrm{km}\mathrm{s}^{-1}\left(\frac{\mu}{\mathrm{mas}}\mathrm{yr}^{-1}\right)\left(\frac{d}{\mathrm{kpc}}\right),$ (26)
@@
Any star with a large $V_{\mathrm{Tan}}$ is likely to have a large $V_{\mathrm{Toomre}}$ and can be assumed to be kinematically a Halo star. However, the selection on $V_{\mathrm{Tan}}$ neglects Halo stars that distinguish themselves from the Disc through high $v_{\mathrm{los}}$. The halo selection cut typically used is $V_{\mathrm{Tan}}>200\,\text{km}\hskip 1.00006pt\text{s}^{-1}$ (Gaia Collaboration et al., 2018; Mikkola et al., 2023), which selects a $V_{\mathrm{Toomre}}$ defined Halo with $83\%$ purity and $68\%$ completeness. A stricter cut improves purity but quickly reduces the completeness, as shown in the middle panel of Fig. 16. $V_{\mathrm{Tan}}>300\,\text{km}\hskip 1.00006pt\text{s}^{-1}$ gives a sample of greater than $99\%$ purity but only $21\%$ completeness.
@@ 8.2 Improving the method
Our methodology could be developed further in several ways. First, we have restricted ourselves to a sample within $2.5\,\mathrm{kpc}$ to ensure a high-quality, well-sampled representation of the phase-space distribution. If we could extend our methodology to greater distances, we would have a greater number of stars and additional structures not seen within the local solar neighbourhood, for example, the Sagittarius stream. This extension would require a sufficiently large 6D sample of stars at larger distances to model the phase space effectively. Alternatively, we could model the action distributions of the Disc and Halo using analytic distributions (Li & Binney, 2022; Binney & Vasiliev, 2024). These distributions would allow us to find the likely Disc and Halo membership probability for 5D stars at any distance and not suffer from poor sampling in high-energy regions. However, the Halo is arguably composed entirely of substructure (Naidu et al., 2020), the majority material of GES. Any functional distribution would be imperfect and could easily dominate our smaller substructures or cause biases in poorly fit regions of the phase space. Instead, we have chosen to remain as close to the data as possible.
@@ 1. Observational probability of a single star at a single point
First, we must be able to estimate the probability that a given star, with known stellar properties and position is in our data sample. When a star of intrinsic absolute magnitude and colour is observed, its apparent magnitude and colour are affected by extinction, depending on the local dust distribution and its physical position relative to the Sun. We use the dust map of Lallement et al. (2022) and the coefficients of Gaia Collaboration et al. (2018) to apply reddening and de-reddening corrections.
The Gaia selection function is complex and the subject of various studies (Boubert & Everall 2022; Castro-Ginard et al. 2023). In Castro-Ginard et al. (2023), the RVS selection function has been estimated to depend on the sky position and apparent magnitude and colour of the star $\left(l,b,G,G-G_{RP}\right)$. Given these observables, these authors provide a method to estimate the probability that the star would be included in the Gaia RVS catalogue. Numerically, the RVS function information is defined with a physical grid of healpix bins and distances, followed by a grid of colours and magnitudes.
@@ 2. Expected observed fraction of a halo population at a single point
Instead of considering the probability of seeing a single star at a point in space, we must consider the expected fraction observed of a typical Halo stellar population. To do this, we draw a stellar population from an isochrone, using a Chabrier (Chabrier 2003) Initial Mass Function (IMF). We use an alpha-enhanced BaSTI isochrone (Pietrinferni et al. 2021) with a metallicity of $\left[\mathrm{Fe}/\mathrm{H}\right]=-1.1$ and an age of $11$ Gyrs.
@@ Appendix B Density estimation
$\rho\left(\sqrt{J_{R}+\delta_{J}},\sqrt{J_{z}+\delta_{J}},L_{z}\right),$
@@
$\rho\left(\tilde{\bm{J}}|\left[\tilde{\bm{J}}_{i},w_{i}\right]\right)={\left(\sigma_{R}\sigma_{Z}\sigma_{\phi}\right)}^{-1}\sum_{i=1}^{N}w_{i}{\lambda_{i}}^{-3}K_{e}\left(\frac{1}{\lambda_{i}}\left|\frac{\tilde{\bm{J}}-\tilde{\bm{J}}_{i}}{\bm{\sigma}}\right|\right),$ (27)
@@
$K_{e}\left(r\right)=\begin{cases}\frac{d+2}{2V_{d}}\left(1-r^{2}\right)&r^{2}<1\\ 0&\text{otherwise},\end{cases}$ (28)
@@
$\sigma_{j}=\frac{P_{100-\omega\%}\left(\tilde{J}_{j}\right)-P_{\omega\%}\left(\tilde{J}_{j}\right)}{\log\left(N\right)},$ (29)
@@
$\lambda_{i}={\left(\frac{\hat{\rho}\left(\tilde{\bm{J}}_{i}\right)}{g}\right)}^{-\frac{1}{3}},$ (30)
@@ Appendix C Disc disequillibirum
Instead, we look at the dynamics within the Disc using the cylindrical components of velocity $\left(v_{R},v_{\phi},v_{Z}\right)$ (using the true RVS $v_{\mathrm{los}}$ values). We consider the deviation from expected equilibrium values to quantify disequilibrium in velocity. If the Disc is in phase-mixed equilibrium in every bin, we expect even numbers of stars moving up and down through the disc and in and out towards the centre of the Galaxy. Defining $N_{+}$ and $N_{-}$ as the number of stars with a positive and negative velocity sign, respectively, we can estimate the statistical significance of any imbalance by modelling the distribution as a binomial, which can be approximated by a normal distribution. The statistical likelihood of getting $\left(N_{+},N_{-}\right)$ in units of sigma is then $\chi$:
$\chi=\frac{1}{\sqrt{N}}\left(N_{+}-N_{-}\right),$ (31)
@@
We plot these $\chi$ values for each of the velocity components in the bottom panels. They show strong, clear evidence for dynamical disequilibrium in Disc stars. The left panel showing the tangential value has very good agreement with the local spiral arm structure, as can be seen in Poggio et al. (2021) Fig. 5. The middle radial panel is consistent with Vislosky et al. (2024), and at the solar position, we are at the edge of a structure that is moving radially outward. The final panel of vertical motion shows a slight structure at a larger radius, away from the centre. There is more structure when comparing above and below the disc plane, which is likely linked to the Gaia phase space spiral (Antoja et al. 2018; Hunt et al. 2022).